Elsa has 12 more pebbles than Ana has. Ana has 6 more pebbles than Olaf has. If Elsa has twice as many
pebbles as Olaf, how many pebbles does each of them have?

Answer by ikleyn(52778)

.
Elsa has 12 more pebbles than Ana has. Ana has 6 more pebbles than Olaf has. If Elsa has twice as many
pebbles as Olaf, how many pebbles does each of them have?
~~~~~~~~~~~~~~~

Olaf has x pebbles

Elsa has (2x) pebbles

Ana has (x+6) pebbles.


We are given

       Elsa = Ana + 12,   

or

       2x = (x+6) + 12


Simplify and find x


       2x = x + 18

       2x - x = 18

          x   = 18


ANSWER.  Olaf has 18 pebbles;  Elsa has 2*18 = 36 pebbles;  Ana has 18+6 = 24 pebbles.

Solved.